OpenStudy (anonymous):

I will be a fan and give medal Find the value of term a14 in the sequence. 7, 2, –3, –8, –13, . . . –58 –53 –63 –27

OpenStudy (godlovesme):

Hold on lemme see if I can find the formula for the sequence..

OpenStudy (gorv):

first term is 7

OpenStudy (gorv):

2-7=-5 that is 5 is subtracted from 2nd term

OpenStudy (gorv):

-3-2=-5 so the common difference =-5 first term =7 and u have to find the 14th term that is n=14 now use the formula

OpenStudy (gorv):

\[t_n=a+(n-1)*d\]

OpenStudy (gorv):

we need to find \[t_{14}\]

OpenStudy (gorv):

a=7 n=14 d=-5

OpenStudy (anonymous):

I got c, is that correct? @gorv

OpenStudy (gorv):

nope

OpenStudy (anonymous):

what would i do with the t14 then

OpenStudy (gorv):

\[t_{14}=7+(14-1)*(-5)\]

OpenStudy (anonymous):

A

OpenStudy (gorv):

yep :)